In a double-slit experiment it is found that blue light of wavelength 467 nm gives a second-order maximum at a certain location on the screen. What wavelength of visible light would have a minimum at the same location?
Answer #1
Given that
The wavelength of the blue light(λb) = 467nm
For the maxima of second order for blue light
We have dsinθ = mλb
               dsinθ= 934nm
For a minimum interference of the other light
               dsinθ = (m+1/2)λ
                934nm =(m+1/2)λ
Where m =0,1,2,3................
For the first 3 values of m we get
If m =0 then 934nm = (1/2)λ then λ = 1868nm
If m=1 then 934nm = (3/2)λ then λ =623nm
If m=2 then 934nm = (5/2)λ then λ = 374nm
Therefore the only wavelength in the visible light is (λ)=623nm